# BehAV: Behavioral Rule Guided Autonomy Using VLMs for Robot Navigation in Outdoor Scenes

### v1

@@ Abstract
We present BehAV, a novel approach for autonomous robot navigation in outdoor scenes guided by human instructions and leveraging Vision Language Models (VLMs). Our method interprets human commands using a Large Language Model (LLM), and categorizes the instructions into navigation and behavioral guidelines. Navigation guidelines consist of directional commands (e.g., “move forward until”) and associated landmarks (e.g., “the building with blue windows”), while behavioral guidelines encompass regulatory actions (e.g., “stay on”) and their corresponding objects (e.g., “pavements”). We use VLMs for their zero-shot scene understanding capabilities to estimate landmark locations from RGB images for robot navigation. Further, we introduce a novel scene representation that utilizes VLMs to ground behavioral rules into a behavioral cost map. This cost map encodes the presence of behavioral objects within the scene and assigns costs based on their regulatory actions. The behavioral cost map is integrated with a LiDAR-based occupancy map for navigation. To navigate outdoor scenes while adhering to the instructed behaviors, we present an unconstrained Model Predictive Control (MPC)-based planner that prioritizes both reaching landmarks and following behavioral guidelines. We evaluate the performance of BehAV on a quadruped robot across diverse real-world scenarios, demonstrating a 22.49% improvement in alignment with human-teleoperated actions, as measured by Fréchet distance, and achieving a 40% higher navigation success rate compared to state-of-the-art methods.
@@ IV BehAV:Behavioral Rule Guided Autonomy Using VLMs
In this section, we outline the methodology of our approach. BehAV is structured into four key components: 1. Human Instruction Decomposition; 2. Behavioral Cost Map Generation; 3. Visual Landmark Estimation; 4. Behavior-Aware Planning. Each of these components is discussed in detail below. Overall system architecture is in Fig. 2.
Figure 3: Sample behavioral cost maps generated by BehAV for different instruction sets: (a)”Follow the sidewalk, stay away from grass, and avoid cyclists”; (b)”Stay on sand, stay away from grass, and avoid water puddles”; (c)”Stay on sidewalk, follow the crosswalk and stop for stop hand gesture”; (d)”Stay on concrete, avoid grass and stop for stop sign”. The cost maps use Jet colormap as shown on the color bar.
@@ IV-D Visual Landmark Estimation
$\text{VLM}:\{I_{\text{marked}},\ell_{\text{frontier}}\}\to G_{img}=[x_{img},y_{img}],$ (8)
@@ IV-E1 Handling Extremely Undesirable Actions
When behavioral objects lie beyond the predicted trajectory horizon used to optimize $J(q_{z})$, the planner cannot generate corresponding actions, as these objects are excluded from the optimization. This becomes critical when highly undesirable actions associated with objects beyond the image horizon (e.g., a stop sign) are ignored. To address this, the planner checks for behavioral objects $l_{i}^{\text{behav}}$ linked to undesirable actions $a_{i}^{\text{behav}}\in\mathcal{A}^{\text{behav}}$. Upon detection, the planner constrains the maximum velocity in the optimizer based on the undesirability of the action. This results in reduced velocities according to the behavioral action cost $p_{i}^{\text{desirable}}$, ensuring safer navigation even when critical objects are beyond the trajectory horizon.
@@ V-A Implementation
For real-time deployment and inference, we use the Ghost Vision 60 robot from Ghost Robotics equipped with an OS1-32 LiDAR, L515 Realsense Camera, an onboard Intel NUC 11, which includes an Intel i7 CPU and an NVIDIA RTX 2060 GPU. Trajectory optimizer is implemented using the NLopt Python package. Code implementation details with default parameters are available on the project webpage.

### v2

@@ Abstract
We present BehAV, a novel approach for autonomous robot navigation in outdoor scenes guided by human instructions and leveraging Vision Language Models (VLMs). Our method interprets human commands using a Large Language Model (LLM), and categorizes the instructions into navigation and behavioral guidelines. Navigation guidelines consist of directional commands (e.g., “move forward until”) and associated landmarks (e.g., “the building with blue windows”), while behavioral guidelines encompass regulatory actions (e.g., “stay on”) and their corresponding objects (e.g., “pavements”). We use VLMs for their zero-shot scene understanding capabilities to estimate landmark locations from RGB images for robot navigation. Further, we introduce a novel scene representation that utilizes VLMs to ground behavioral rules into a behavioral cost map. This cost map encodes the presence of behavioral objects within the scene and assigns costs based on their regulatory actions. The behavioral cost map is integrated with a LiDAR-based occupancy map for navigation. To navigate outdoor scenes while adhering to the instructed behaviors, we present an unconstrained Model Predictive Control (MPC)-based planner that prioritizes both reaching landmarks and following behavioral guidelines. We evaluate the performance of BehAV on a quadruped robot across diverse real-world scenarios, demonstrating a 22.49% improvement in alignment with human-teleoperated actions, as measured by Fréchet distance, and achieving a 40% higher navigation success rate compared to state-of-the-art methods. Code and video are available at http://gamma.umd.edu/behav/.
@@ IV BehAV:Behavioral Rule Guided Autonomy Using VLMs
In this section, we outline the methodology of our approach. BehAV is structured into four key components: 1. Human Instruction Decomposition; 2. Behavioral Cost Map Generation; 3. Visual Landmark Estimation; 4. Behavior-Aware Planning. Each of these components is discussed in detail below. The overall system architecture is in Fig. 2 and the algorithm is summarized in Algorithm 1.
Figure 3: Cost maps generated by BehAV for diverse instructions: (a)“Follow the sidewalk, stay away from grass, and avoid cyclists”; (b)“Stay on the sand, stay away from grass, and avoid water puddles”; (c)“Stay on the sidewalk, follow the crosswalk and stop for stop hand gesture”; (d)“Stay on concrete, avoid grass and stop for stop sign”; (e)“Stay on tiles, and use caution to follow stairs”; (f)“Follow the concrete, stay away from grass, and stop for people wearing red shirts”; (g)“Stay on concrete, stay away from grass and yield to people wearing black shirts”; (h)“Stay on concrete, stay away from the grass”; (i)“Stay on grass, stay away from the concrete”. The color map is shown on the right side.
@@ IV-D Visual Landmark Estimation
$\text{VLM}:\{I_{\text{RGB}},\ell_{\text{frontier}},l^{\text{nav}}_{i}\}\to G_{img}=[x_{img},y_{img}],$ (8)
@@ IV-E Behavior-Aware Planning
The optimal trajectory parameterization $z^{*}$ obtained by optimizing the cost function $J(q_{z})$ is used to calculate linear and angular velocities $(v,\omega)_{t}$ for the current time instant $t$ using the control law in Eq. 1.
@@ IV-E1 Handling Extremely Undesirable Actions
When behavioral objects lie beyond the predicted trajectory horizon used to optimize $J(q_{z})$, the planner cannot generate corresponding actions, as these objects are excluded from the optimization. This becomes critical when highly undesirable actions associated with objects beyond the image horizon (e.g., a stop sign) are ignored. To address this, the planner checks for behavioral objects $l_{i}^{\text{behav}}$ linked to undesirable actions $a_{i}^{\text{behav}}\in\mathcal{A}^{\text{behav}}$. Upon detection, the planner constrains the maximum velocity in the optimizer based on the undesirability of the action.
Let, $\mathcal{B}_{\text{lower}}=\{r_{\text{min}},\theta_{\text{min}},\delta_{\text{min}},v_{\text{min}}\}$ and $\mathcal{B}_{\text{upper}}=\{r_{\text{max}},\theta_{\text{max}},\delta_{\text{max}},v_{\text{max}}\}$ be the lower and upper bounds of $z$ for the trajectory optimizer, we constrain the $v_{\text{max}}$ in the upper bound as, $v_{\text{max}}^{\prime}=(1-\text{max}(\mathcal{C}_{behav})).v_{\text{max}}$ such that the maximum achievable velocity for the robot is significantly lower or zero in the presence of objects associated with extremely undesirable actions (i.e., $\text{max}(\mathcal{C}_{behav})\geq c_{\text{th}}$, where $c_{\text{th}}$ is a threshold).
This results in reduced velocities according to the behavioral action cost $p_{i}^{\text{desirable}}$, ensuring safer navigation even when critical objects are beyond the trajectory horizon.
@@ IV-E2 Gait Switching for Stability
1:  Input: $\ell_{\text{instr}},\ell_{\text{prompt}},\ell_{\text{action}},\ell_{\text{frontier}},I_{RGB,t},P_{t}^{lidar},odom$
2:  Output: $(v,\omega)_{t},gait_{\text{caution}}$
3:  Initialize : $k_{1},k_{2},T,w_{\text{goal}},w_{\text{obs}},w_{\text{behav}},d_{\text{safe}},d_{\text{th}},\lambda,r_{\text{max}},$ $r_{\text{min}},\theta_{\text{max}},\theta_{\text{min}},\delta_{\text{max}},\delta_{\text{min}},v_{\text{max}},v_{\text{min}},c_{\text{th}},gait_{\text{caution}}$
4:  $\text{LLM}:\{\ell_{\text{instr}},\ell_{\text{prompt}}\}\to\{\mathcal{A}^{\text{nav}},\mathcal{L}^{\text{nav}},\mathcal{A}^{\text{behav}},\mathcal{L}^{\text{behav}}\}$
5:  $\text{LLM}:\{\mathcal{A}^{\text{behav}},\ell_{\text{action}}\}\to\mathcal{P}^{\text{desirable}}$
6:  $\text{VLM}:\{I_{\text{RGB}},\ell_{\text{frontier}},l^{\text{nav}}\}\to G_{img}=[x_{img},y_{img}]$
7:  $G_{odom}=\mathcal{T}_{img}^{\mathcal{O}}G_{img}$
8:  $d_{\text{goal}}=\|G_{odom}-odom\|_{2}$
9:  $\mathcal{B}_{\text{lower}}=\{r_{\text{min}},\theta_{\text{min}},\delta_{\text{min}},v_{\text{min}}\}$
10:  $\mathcal{B}_{\text{upper}}=\{r_{\text{max}},\theta_{\text{max}},\delta_{\text{max}},v_{\text{max}}\}$
11:  while $d_{\text{goal}}\geq d_{\text{th}}$ do
12:     $d_{\text{goal}}=\|G_{odom}-odom\|_{2}$
13:     Obtain $S_{i}\forall l_{i}^{\text{behav}}\in\mathcal{L}^{\text{behav}}$ from CLIPSeg [34] (Eq. 5)
14:     Calculate $\mathcal{C}_{behav}$ using Eq. 6 and 7.
15:     if $\text{max}(\mathcal{C}_{behav})\geq c_{\text{th}}$ then
16:        $v_{\text{max}}^{\prime}=(1-\text{max}(\mathcal{C}_{behav})).v_{\text{max}}$
17:        $\mathcal{B}_{\text{upper}}=\{r_{\text{max}},\theta_{\text{max}},\delta_{\text{max}},v_{\text{max}}^{\prime}\}$
18:     end if
19:     Optimize $J(q_{z})$ for $z\in[\mathcal{B}_{\text{lower}},\mathcal{B}_{\text{upper}}]$
20:     $z^{*}=\arg\min_{z}J(q_{z})$
21:     Calculate $(v,\omega)$ for $z^{*}$ using Eq. 1
[…]
23:        $gait_{\text{caution}}=True$
24:     end if
25:     return  $(v,\omega),gait_{\text{caution}}$
26:  end while
Algorithm 1 Outdoor Navigation using BehAV
@@ V-A Implementation
For real-time deployment and inference, we use the Ghost Vision 60 robot from Ghost Robotics equipped with an OS1-32 LiDAR, L515 Realsense Camera, an onboard Intel NUC 11, which includes an Intel i7 CPU and an NVIDIA RTX 2060 GPU. Trajectory optimizer is implemented using the NLopt Python package. We use a combination of global and location optimization to achieve the desired trajectory optimization. Code implementation details with default parameters are available on the project webpage.